OpenStudy (anonymous):

for what real numbers is x is x^2-6x+9 negative? A. -33 C. x=-3 or x=3 D. 0< x<6 E. for no real number x

OpenStudy (anonymous):

\[x^2-6x+9=\left( x-3 \right)^2 \ge 0\]

OpenStudy (anonymous):

@surjithayer how does it go to (x-3)^2> 0

OpenStudy (anonymous):

\[x^2-2*x*3+3^2=\left( x-3 \right)^2\] square of a real number is always positive.

OpenStudy (anonymous):

\[\left( a-b \right)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

is b?

OpenStudy (chealyn98):

B

Directrix (directrix):

B is not correct. And, a reason for that choice should be given.

OpenStudy (anonymous):

what do you mean by that @Directrix

Directrix (directrix):

Two Things: The answer to this question is not B. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (rizags):

First Factor it to:\[(x-3)(x-3)\] then from here we can analyze the signs of the parts: 

OpenStudy (rizags):

Now guess what happens when we multiply those one either side of 3

OpenStudy (rizags):

those one either side of 3. Take a point, and multiply the signs on either side and see what you get!

Directrix (directrix):

for what real numbers is x is x^2-6x+9 **negative**?

OpenStudy (anonymous):

so we multiply it by its self?

Directrix (directrix):

Name a value of x for which (x - 3) ^2 is negative.

OpenStudy (anonymous):

whats a real; number?

Directrix (directrix):

Name any number that makes this negative: (x - 3) ^2

Directrix (directrix):

Read about real numbers here: http://www.mathsisfun.com/numbers/real-numbers.html

OpenStudy (anonymous):

another (x-3)

Directrix (directrix):

If you square any real number, the result will never be negative.

OpenStudy (anonymous):

true?

Directrix (directrix):

Here is the incorrect option B: B. x<-3 or x>3 It says that if I pick any number greater than 3, then (x - 3) ^2 will be negative. Let us test that. (x - 3) ^2 x = 4 (4 - 3) ^2 = 1^2 = 1 which is POSITIVE and not negative. Option B is wrong.

Directrix (directrix):

If you square a real number you will NEVER get a negative product.

OpenStudy (anonymous):

Ok thanks